Solve the formula for the given letter: A=1/2bh for h
Answer by ptaylor(2198)

Solve the formula for the given letter:
A=1/2bh for h
This looks like 1 over 2bh but I think that it should be 1/2 times (bh) which is the area of a triangle. That's what I will assume:
A=(1/2)(bh) multiply both sides by 2
2A=bh divide both sides by b
2A/b=bh/b and
h=2A/b
Hope this helps----ptaylor